OpenStudy (fools101):

Find the zeros of the following equations/functions y = x2 + 2x + 3

OpenStudy (fools101):

please help!

jhonyy9 (jhonyy9):

x_1,_2=(-2+/- sqrt(4-12))/2 =(-2+/- sqrt(-8))/2=2(-1+/-isqrt2)/2 = = -1+/- isqrt2

OpenStudy (anonymous):

Possible intermediate steps: x^2+2 x+3 = 0 Subtract 3 from both sides: x^2+2 x = -3 Add 1 to both sides: x^2+2 x+1 = -2 Factor the left hand side: (x+1)^2 = -2 Take the square root of both sides: sqrt(x+1) = i sqrt(2) Eliminate the absolute value: x+1 = -i sqrt(2) or x+1 = i sqrt(2) Subtract 1 from both sides: x = -1-i sqrt(2) or x+1 = i sqrt(2) Subtract 1 from both sides: x = -1-i sqrt(2) or x = -1+i sqrt(2)

OpenStudy (fools101):

Thank u!!!